The sum of three consecutive odd integers is 189. What are the integers?
Answer by lenny460(1073)

The sum of three consecutive odd integers is 189. What are the integers?

Let the three consecutive odd integers be:
n + 1, n + 3 and n + 5
Therefore:
n + 1 + n + 3 + n + 5 = 189
3n + 9 = 189
3n = 180
n = 60
n + 1 = 61
n + 3 = 63 and
n + 5 = 65

Answer:
The three consecutive odd integers are:
61, 63 and 65




Lennox Obuong
Algebra Tutor